Your research department estimates that the supply function for televisions is given by:

QXS = 5,000 + 5PX -10PR – 2PW

   When PX is $500, PR is $100, PW is $2500, how many television sets are produced?
Answer #1

The supply function lets us estimate the quantity supplied of a good when the market conditions are given.

In this question , supply function is :

QXS = 5000 + 5PX - 10PR - 2PW

Putting the values given in the equation, we get

QXS = 5000+(5*500)-(10*100)-(2*2500)

QXS = 5000+2500-1000-5000

QXS = 1500

So , with these values of the variables , the number of television sets that are produced are 1500.